What is the equation of the line that contains (13,-11) and has a slope of 2.?
Answer by mananth(16946)

m= 2

Plug value of the slope and point ( 13 , -11 ) in
Y = m x + b
-11.00 = 26 + b
b= -11 - 26
b= -37
So the equation will be
Y = 2 x -37
m.ananth@hotmail.ca